# Many Worlds in Theory Space: A Quantum Origin for the Constants of Nature

### v1

@@ 8.5 Relation to Previous Work
The conceptual unification of the cosmological multiverse with the Many-Worlds Interpretation has been proposed previously, notably by Nomura [Nomura2011] and Susskind/Bousso [BoussoSusskind2012]. However, these approaches typically rely on the semiclassical background of eternal inflation and bubble dynamics.
This work distinguishes itself by providing a fully canonical realization of this unification. Instead of relying on bubble nucleation, we derive the splitting of physical laws directly from the decoherence of the universal wavefunction in the extended configuration space. By utilizing the Wheeler-DeWitt formalism, we show how “theory selection” emerges as a stationary feature of the quantum state $\Psi$, independent of the specific temporal slicing issues associated with the measure problem in eternal inflation.
@@ 9 Conclusion
Finally, this framework makes a specific, falsifiable prediction if it is the correct one: there will never be a successful, purely mathematical derivation of the Standard Model parameters.** The ultimate theory will yield a probability distribution, not a point solution. This suggests that the laws of physics we observe are not fixed external constraints, but frozen quantum accidents—local fossils of the early universe’s decoherence history.

### v2

@@ 8.4 Relation to Previous Work
The unification of the cosmological multiverse with the Many-Worlds Interpretation has been proposed previously, notably by Carr & Rees[CarrRees1979], Nomura [Nomura2011] and Susskind/Bousso [BoussoSusskind2012] in concept form. This work distinguishes itself by providing a fully canonical realization of this unification. Instead of relying on bubble nucleation, we derive the splitting of physical laws directly from the decoherence of the universal wavefunction in the extended configuration space. By utilizing the Wheeler-DeWitt formalism, we show how “theory selection” emerges as a stationary feature of the quantum state $\Psi$, independent of the specific temporal slicing issues associated with the measure problem in eternal inflation.
@@ 9 Conclusion
It is important to distinguish the role of this framework from that of a fundamental unified theory like String Theory. String Theory provides the microscopic description of the ’Theory Space’ $\mathcal{T}$. It defines the manifold of possible vacua, the allowed gauge groups, and the relationships between moduli and coupling constants. This work provides the selection mechanism: it describes how the quantum state of the universe evolves across this landscape, decoheres, and stabilizes into specific U-sectors. Thus, our Extended MWI does not replace String Theory; rather, it completes it by providing the dynamical context in which the landscape of possibilities is realized
@@
Finally, this framework allows comparisons to be made between competing quantum cosmology theories and makes a specific, falsifiable prediction if it is the correct one: there will never be a successful, purely mathematical derivation of the Standard Model parameters. The ultimate theory will yield a probability distribution, not a point solution. This suggests that the laws of physics we observe are not fixed external constraints, but frozen quantum accidents; local fossils of the early universe’s decoherence history.